OpenStudy (anonymous):

Consider an exponential random variable Expμ , with μ=4.5 . find a so that P(Expμ∈(0,a])=0.65 .

OpenStudy (anonymous):

does this mean is \[6.5 = \int\limits_{0}^{a} \frac{ 1 }{ 4.5 } e^{-{\frac{ x }{ 4.5 }}} dx \] ?

OpenStudy (amitrashok):

6

OpenStudy (anonymous):

no is wrong calculus